Determine the net heat radiated by the black body disc of mass 6g and radius 4cm at 400C,initially kept at 300C.
1
Expert's answer
2020-12-21T04:02:39-0500

According to Stefan-Boltzmann's the loose rate

"\frac{\partial Q_{out}}{\partial t}=\pi r^{2}\sigma (T_{sur}^{4}-T^{4})\\

 \sigma = 5.67\times10^{-8}\times\frac{W}{m^{2}K^{4}}\\
\frac{22}{7}\times(0.04)^{2}\times5.67\times 10^{-8}(400^{4}-300^{4})\\
=4.9896J"